[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Apr 20, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Apr 20 19:28:17 UTC 2017


HSU Shift Log for Thu Apr 20 2017 15:28:10 GMT-0400 (EDT)

Shift: H

Forecaster: L Avila

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: P-Surge was run Hot line test was done.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170420/e8fb2775/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list